HUBBLE SPACE TELESCOPE DAILY REPORT #1177 PERIOD COVERED: 0700 EST Daily Status Report as of 192/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Proposal HRS 5319 The Faint Object Spectrograph was used used to observe star Fiege 46, an O-type star known to have peculiar chemical abundances. Ten observations were received and displayed in the OSS. They were reported to look okay. The target was found on the sixth map of the 6X6 spiral search target acquisition. No anomalies were noted. 1.2 Completed Proposal 5641 FGS-3 was used to make astrometric observations of asteroid PAUL3 to study its positional accuracy. The first set of observations occurred during LOS. After the telemetry was returned, no problems were seen, until the end when the stop flag was set. No problems were seen with the next two observations. On subsequent observations, it was reported that there were no successful observations on target PAUL3, but the observations of the reference object appeared to be successful. The scan limit was reported exceeded and the stop flag was set. 1.3 Completed Proposal 5377 The Faint Object Camera (f/96) was used to make an internal flat calibration image and an ultraviolet observations of the known X-Ray pulsar PSR1929+10 to see if its ultraviolet flux matches with the model of neutron star atmospheres. The observations were executed as scheduled, and the observations remain on Science Tape Recorder #3. The data will be analyzed later. No anomalies were noted. 1.4 Completed Proposal 5397 The WF/PC-2 was used to make images of galaxy NGC-7331 to be used in measuring the periods of Cepheid variable stars within the galaxies. The observations were executed as scheduled, and the observations remain on Science Tape Recorder #3. The data will be analyzed later. No anomalies were noted. 1.5 Completed Proposal 5450 The WF/PC-2 was used to make observations of Quasar QSO-114118 to study the properties of its host galaxy and surrounding environment. The observations were executed as scheduled. The data remains on Science Tape Recorder #3, and will be analyzed later. 1.6 Completed Proposal 5562 The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Proposal 5370 The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations as part of a medium-deep sky survey. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Proposal 3856 FGS-3 were used to make parallax astrometric observations of star G166-37. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Proposal 5442 The WF/PC-2 was used to make observations of the gravitionally lensed quasi-stellar system Q1413+117 (aka. the Clover Leaf). The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled :24 Successful :22 There were two FHST failures during safemode recovery due to attitude errors. 2.3 Significant HSTARs: None 2.4 Operations Notes: Normal science observations was successfully resumed on Saturday evening at 6:00 P.M. All observatory subsystems performing nominally. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END- .